# UniVST: A Unified Framework for Training-free Localized Video Style Transfer

### v3

@@ 4.1 Point-Matching Mask Propagation
To preserve the primary style during the style transfer process, a potent technique is to apply a mask to blend the latent variables at several steps. This can be expressed as:
$\mathcal{Z}_{t}=M\cdot Z_{t}+(1-M)\cdot\mathcal{Z}_{t},$ (5)
where $\mathcal{Z}_{t}$ and $Z_{t}$ denote the latent variables in the editing and content branch, respectively. However, in contrast to images, generating masks $M$ for each video frame is laborious. Users typically prefer to control the video’s main style by supplying a mask for only the initial frame. Extending the user-provided mask across all frames efficiently is necessary. Thus, we introduce point-matching mask propagation below and details are provided in Algorithm 1.
Figure 3: Comparison of accuracy and inference time under different mask propagation strategies. The introduction of anchor frames with the downsampling strategy significantly mitigates the accuracy degradation in subsequent frame propagation and also reduces inference time.
@@
Point Matching. Inspired from DIFT [32], the three upsampling blocks in the UNet are replete with rich semantic information. Therefore, we can retain the feature map $\{F_{t_{0}}^{i}\}_{i=1}^{N}$ from upsampling block-2 at a given step $t_{0}$ during DDIM inversion in the source video, and use them to calculate the pixel correspondences between frames. Specifically, to pinpoint the pixel in the $j$-th frame that most closely resembles a given point $p_{i}$ in the $i$-th frame, we calculate the cosine similarity between the corresponding point in the feature map $F_{t_{0}}^{i}$ and all points in the feature map $F_{t_{0}}^{j}$ as:
$p_{j}={\underset{p_{j}\in F_{t_{0}}^{j}}{\arg\min}}\ CosSim\big{(}p_{i},p_{j}\big{)},\quad p_{i}\in F_{t_{0}}^{i}.$ (6)
Utilizing this principle, we can propagate the mask from the first frame to all subsequent frames using the mask propagation strategy formally introduced in the following.
@@
Mask Propagation. Given an initial object mask for the first frame from the user, we use $k$-NN [33] to propagate this mask to subsequent frames. For each point in the $i$-th frame, we find the $k$ most similar points in the $(i-1)$-th frame using point matching. If the majority of these $k$ points are in the foreground, the corresponding point in the $i$-th frame is also classified as foreground; otherwise, it is background.
Figure 4: (a) Original video frame; (b) Key-value replacement [17]; (c) Key-value AdaIN; (d) Key-value AdaIN with gradual attention shift.
However, as shown in Fig. 3, using a naive autoregressive approach for mask propagation can lead to error accumulation, significantly decreasing accuracy for subsequent frames. To address this, we design the anchor frames mechanism that incorporates information from the first frame and the previous $n$ frames.
Although introducing such a mechanism can improve the propagation accuracy of subsequent frames, it requires more similarity calculations, potentially reducing efficiency. To address this issue, we apply random downsampling at a rate of $r$ to both foreground and background regions of the anchor frames to minimize costs. The sampling rates are then adjusted to reflect the proportions in the foreground and background. As illustrated in Fig. 3, the combination of anchor frames and random downsampling maintains final mask propagation accuracy while improving efficiency. More comprehensive information and in-depth analysis can be available in the experimental section.
4.2 AdaIN-Guided Video Style Transfer
By using DDIM inversion [34] to the original video $\{I^{i}\}_{i=1}^{N}$ and the reference style image $I^{s}$, we obtain their respective noise latents $Z_{t}$ and $Z_{t}^{s}$ ($t=1\rightarrow T$). Then, we establish the edited latents $\mathcal{Z}_{t}=Z_{t}$, and integrate a three-branch architecture with latent-shift and attention-shift mechanisms to accomplish video style transfer.
AdaIN-Guided Latent Shift. We have discovered that applying the AdaIN [17] to the initial noise significantly enhances the transferred color style. Therefore, we extend the application of AdaIN to several steps during the denoising process. Within the interval $t\in[\tau_{0},\tau_{1}]$, We consider applying it to the latent $\mathcal{Z}_{t}$ of the editing branch, shifting its mean and variance to the latent $Z_{t}^{s}$ of the style branch:
$\mathcal{Z}_{t}$ $=\text{AdaIN}(\mathcal{Z}_{t},Z_{t}^{s}),$ (7)
@@ 4.2 AdaIN-Guided Video Style Transfer
where $\mu(\cdot)$ and $\sigma(\cdot)$ denoting mean and standard deviation.
AdaIN-Guided Attention Shift. To prevent the direct replacement of self-attention from causing content damage, [17] blended the original video’s query $Q_{t}$ with the edited query $\mathcal{Q}_{t}$ as:
$\begin{split}\tilde{\mathcal{Q}}_{t}&=\gamma\cdot\mathcal{Q}_{t}+(1-\gamma)\cdot Q_{t},\\ &\tilde{\mathcal{K}}_{t}=K_{t}^{s},\;\tilde{\mathcal{V}}_{t}=V_{t}^{s},\end{split}$ (9)
where $\gamma$ controls the balance between content and style. However, as observed in Fig. 4(b), this approach still lead to a loss of localized details in video stylization. To mitigate the loss of details caused by direct replacement, we propose to shift the mean and variance of the distributions of $\mathcal{K}_{t}$ and $\mathcal{V}_{t}$ using AdaIN:
$\tilde{\mathcal{K}}_{t}=\text{AdaIN}(\mathcal{K}_{t},K_{t}^{s}),\;\tilde{\mathcal{V}}_{t}=\text{AdaIN}(\mathcal{V}_{t},V_{t}^{s}).$ (10)
Though this shift preserves localized details, as indicated in Fig. 4(c), it causes content blurring and artifacts. Upon analysis, we find that the early denoising stages contain less information, making key and value replacements inappropriate. Thus, within the interval $t\in[\tau_{2},\tau_{3}]$, we gradually decrease the impact of the attention shift by combining the key and value replacements with the query blending:
@@
$\beta_{t}$ $=\frac{\beta_{\tau_{3}}-\beta_{\tau_{2}}}{\tau_{3}-\tau_{2}}\cdot(t-\tau_{2})+\beta_{\tau_{2}}.$
By combing key-value AdaIN with gradual attention shift in Eq. (11) and query blending in Eq. (9), the stylization results achieve an effective balance between content fidelity and style richness, as shown in Fig. 4(d) and Table VII.
@@ 4.3 Sliding-Window Consistent Smoothing
We address the common issues of flicker and artifacts in video stylization by focusing on temporal inconsistency. Following [1, 2], we utilize optical flow with specialized sliding window smoothing.
Optical Flow Warping. Optical flow estimation is a pivotal technique, often used for object tracking [35, 36]. It predicts the motion direction and velocity of pixels by analyzing changes between consecutive frames. Using models like RAFT [37], optical flow information can be predicted from one frame to the next, allowing for the reconstruction of the initial frame. In cases where occlusion occurs, bidirectional optical flow can estimate the mask and fill in occluded regions with the original image. This process, termed “Warp,” is key to our following discussions.
Sliding Window Smoothing. We apply the warp operation during denoising process to achieve local smoothing. At time step $t$, we calculate the latent representation $\mathcal{Z}_{t\rightarrow 0}$ from the predicted noise $\varepsilon_{t}$ according to Eq. (2). $\mathcal{Z}_{t\rightarrow 0}$ is then decoded by the decoder $\mathcal{D}$ into a pixel representation $\{P_{t}^{i}\}_{i=1}^{N}$. For each decoded frame $P_{t}^{i}$, we apply a sliding window of size $2m$ to warp its neighboring frames and smooth them to update the original pixel:
@@
The updated values from the current window will be incorporated into the calculations for the next window, as shown in Fig. 2. Subsequently, the VAE’s encoder $\mathcal{E}$ re-encodes the smoothed $\{\bar{P}_{t}^{i}\}_{i=1}^{N}$ into latent representations $\bar{\mathcal{Z}}_{t\rightarrow 0}$, which is an updated version of $\mathcal{Z}_{t\rightarrow 0}$. Next, the predicted noise $\varepsilon_{t}$ is adjusted using the inversion of Eq. (2):
@@
Finally, the latent representation is updated in accordance with the DDIM schedule in Eq. (1):
@@
It is important to note that we do not adopt a global smoothing strategy because extensive optical flow predictions would introduce significant computational overhead. Instead, by employing a sliding local window scheme, we achieved a balance between temporal consistency and efficiency. And this scheme is only performed within the interval $t\in[\tau_{4},\tau_{5}]$.
TABLE I: Quantitative comparison of existing methods across various metrics, superscript * indicates non-diffusion methods. The bold values represent the best performance and the underscore stresses the second best.
Datasets Methods Overall Transfer Performance Style Preservation Consistency Efficiency
@@
DAVTG-Laion Diffutoon [12] 38.610 23.700 0.563 0.959 0.953 0.979 44.263
@@ 5.1 Experimental Settings
Our method is based on the Stable Diffusion V1.5 model, as referenced in the literature [29]. The hyperparameters are configured as follows: $\tau_{0}=0.1T$, $\tau_{1}=0.2T$, $\tau_{2}=0.4T$, $\tau_{3}=1.0T$, $\tau_{4}=0.5T$, $\tau_{5}=0.6T$, $t_{0}=0.4T$, with the total time period $T$ set to 50. Additionally, we have $\gamma=0.35$, $\beta_{\tau_{2}}=0.1$, $\beta_{\tau_{3}}=0.9$, $r=0.3$, $k=15$, $m=2$, and $n=9$. Videos are resized to a resolution of 512 by 512 pixels and are processed in batches consisting of 16 frames each. For optical flow estimation, we employ the RAFT model [37]. Our approach does not require training or fine-tuning and can be executed on a single RTX 3090 GPU, making it both efficient and practical for deployment.
@@
Datasets. Building upon the foundational work referenced in [17, 26], we have meticulously selected style images from two renowned databases: WikiArt [40] and Laion-Aesthetics-6.5+ [41], each contributing a collection of 20 distinct images. For the content videos, we have procured a dataset comprising 50 videos from the DAVIS2016 [42] and TGVE [43] datasets, which we have amalgamated and designated as DAVTG. Through the integration of these datasets, we have successfully established two comprehensive subsets: DAVTG-WikiArt, which contains 1,000 curated pairs, and DAVTG-Laion, which also features 1,000 paired entries.

### v4

@@ 4.1 Point-Matching Mask Propagation
“A towering building begins with a solid foundation.”
As the proverb clearly illustrates, the foundation is crucial for any significant endeavor. In the context of the localized video style transfer, the first essential step is to obtain precise and accurate masks for the region of interest specified by the users. These masks essentially form the cornerstone for enabling fine-grained control over specific regions during the subsequent stylization process.
Compared to image style transfer, the video domain presents a unique challenge: obtaining mask information for each frame. One potential solution is to employ segmentation models, such as SAM [35], which can perform frame-by-frame segmentation based on the specified object. However, as shown in Fig. 3, this approach has notable drawbacks. Specifically, it introduces an additional memory overhead of 7.9 GB and is computationally intensive, which is unfavorable for deploying our UniVST framework.
Given the substantial redundancy between consecutive video frames, we propose leveraging this redundancy to efficiently propagate the initial frame’s mask across all subsequent frames, thereby eliminating the need for additional segmentation models. This approach not only reduces memory usage but also accelerates the overall processing time. Thus, we introduce our point-matching mask propagation strategy, which is outlined in Algorithm 1.
@@
Point-Matching. Inspired by DIFT [36], the three up-sampling blocks in the U-Net are replete with rich semantic information. Therefore, during the DDIM inversion in the original video, we can retain the feature map $\{F_{t_{0}}^{i}\}_{i=1}^{N}$ from the last up-sampling block at a given step $t_{0}$, and use them to calculate the pixel correspondences across frames. Specifically, to pinpoint the pixel in the $j$-th frame that most closely resembles a given point $p_{i}$ in the $i$-th frame, we calculate the cosine similarity between the corresponding point $p^{i}$ in the feature map $F_{t_{0}}^{i}$ and all points $p^{j}$ in the feature map $F_{t_{0}}^{j}$:
$p^{j}={\underset{p^{j}\in F_{t_{0}}^{j}}{\arg\min}}\ \operatorname{CosSim}\big{(}p^{i},p^{j}\big{)},\quad p^{i}\in F_{t_{0}}^{i}.$ (5)
In which, we can extend the initial frame’s mask across subsequent frames by the mask propagation strategy introduced below.
@@
Mask Propagation. Given an initial object mask $\mathcal{M}^{1}$ specified by the users, we use $k$-NN [37] to propagate this mask to subsequent frames. For each point in the $i$-th frame, we identify the $k$ most similar points in the $(i-1)$-th frame based on the point-matching principle described above. If the majority of these $k$ points are in the foreground, the corresponding point in the $i$-th frame is classified as the foreground; otherwise, it is classified as the background. In this way, we can obtain the masks $\{\mathcal{M}^{i}\}_{i=1}^{N}$ for all frames.
However, as shown in Fig. 4, using a naive autoregressive approach for mask propagation will result in error accumulation, which degrades the accuracy of subsequent frames. To mitigate this issue, we design the anchor frames mechanism that incorporates information from the first frame as well as the previous $n$ frames. On the other hand, although the introduction of the anchor frames mechanism improves the propagation accuracy of subsequent frames, it introduces additional similarity calculations, potentially reducing efficiency. To address this issue, we apply random down-sampling at a rate of $r$ to the foreground and background of the anchor frames, with the rate adjusted based on their respective proportions. As illustrated in Fig. 4, the combination of anchor frames and random down-sampling strategies effectively preserves mask propagation accuracy while improving efficiency. More comprehensive analysis will be provided in Sec. 5.4.2.
Figure 4: Comparison of accuracy and inference time under different mask propagation strategies. The accuracy is measured by relevant segmentation metrics: the intersection over union (IoU) and dice coefficients (Dice), which assess the overlap between predicted and ground truth regions. Our strategy effectively balances inference time and accuracy.
4.2 AdaIN-Guided Localized Video Stylization
To achieve localized video stylization, we first separately obtain the initial latent of the original video $\{I^{i}\}_{i=1}^{N}$ and the style image $I^{s}$ using the inversion technique [38], denoted as $Z_{t}$ and $Z_{t}^{s}$ respectively, for $t=1\rightarrow T$. Integrated with our point-matching mask propagation strategy, we can also obtain the mask frame-by-frame efficiently. We then initialize the edited latent $\mathcal{Z}_{T}=Z_{T}$ and introduce a three-branch architecture with content, edited, and style branches, which perform the subsequent denoising process in parallel for $t=T\to 1$. These branches will interact through localized latent blending, AdaIN-guided latent-shift, and AdaIN-guided attention-shift, enabling precise control in subsequent video stylization.
Localized Latent Blending. To ensure that the style within the user-specified region is preserved throughout the denoising process, we utilize the masks $\{\mathcal{M}^{i}\}_{i=1}^{N}$, which is obtained from the previous mask propagation, as auxiliary constraints. Specifically, during each denoising step, we blend the latent of content branch $Z_{t}^{i}$ and the latent of style branch $\mathcal{Z}_{t}^{i}$, using the mask $\mathcal{M}^{i}$ of the $i$-th frame:
$\mathcal{Z}_{t}^{i}=\mathcal{M}^{i}\cdot Z_{t}^{i}+(1-\mathcal{M}^{i})\cdot\mathcal{Z}_{t}^{i},\quad i=1,...,N$ (6)
This blending operation effectively ensures that the style within the user-specified region is seamlessly retained while allowing the rest of the frame to be fully stylized according to the edited branch.
AdaIN-Guided Latent-Shift. AdaIN [24] is primarily used in image processing [17], where it adjusts the mean and standard deviation of an image to control its color style. Inspired by this, we apply the AdaIN technique to integrate the color style of the reference image into the edited video. Specifically, within the interval $t\in[\tau_{0},\tau_{1}]$, we apply AdaIN operation to the latent $\mathcal{Z}_{t}$ of the editing branch, shifting its mean and standard deviation to match those in the latent $Z_{t}^{s}$ of the style branch, as follows:
$\tilde{\mathcal{Z}}_{t}$ $=\text{AdaIN}(\mathcal{Z}_{t},Z_{t}^{s}),$ (7)
@@ 4.2 AdaIN-Guided Localized Video Stylization
where $\mu(\cdot)$ and $\sigma(\cdot)$ denote respective mean and standard deviation.
Figure 5: Comparison of training-free stylization methods: (a) original video frame, (b) key-value replacement [17], (c) key-value AdaIN and (d) AdaIN-guided attention-shift. Our approach mitigates local detail loss.
AdaIN-Guided Attention-Shift. In addition to transferring color tone, we also need to incorporate the overall style of the reference image into the edited video. The previous work StyleID [17] achieves style transfer by modifying the standard self-attention process. Specifically, the query $\mathcal{Q}_{t}$ in the editing branch attends to the key $K_{t}^{s}$ and value $V_{t}^{s}$ of the style branch instead of its own, referred to as K-V replacement. Furthermore, to eliminate content damage caused by direct replacement, [17] blends the original video’s query $Q_{t}$ with the edited query $\mathcal{Q}_{t}$. It is formulated as:
$\begin{split}\tilde{\mathcal{Q}}_{t}&=\gamma\cdot\mathcal{Q}_{t}+(1-\gamma)\cdot Q_{t},\\ &\tilde{\mathcal{K}}_{t}=K_{t}^{s},\;\tilde{\mathcal{V}}_{t}=V_{t}^{s},\\ &\operatorname{Attention}(\tilde{\mathcal{Q}}_{t},\tilde{\mathcal{K}}_{t},\tilde{\mathcal{V}}_{t}),\end{split}$ (9)
where $\gamma$ controls the balance between content and style. However, as observed in Fig. 5(b), K-V replacement still results in slight local artifacts in the stylized videos. To further reduce the local artifacts caused by direct replacement, we apply AdaIN to $\mathcal{K}_{t}$ and $\mathcal{V}_{t}$ in the editing branch, shifting their mean and standard deviation to align with $K^{s}$ and $V^{s}$ in the style branch, which can be described as:
$\tilde{\mathcal{K}}_{t}=\text{AdaIN}(\mathcal{K}_{t},K_{t}^{s}),\quad\tilde{\mathcal{V}}_{t}=\text{AdaIN}(\mathcal{V}_{t},V_{t}^{s}).$ (10)
Though the K-V AdaIN alleviates the artifact phenomenon, as indicated in Fig. 5(c), it introduces new local detail blurring. Upon analysis, we find that early denoising stages have less semantic information, making them suitable for K-V AdaIN, while later stages, containing more semantic information, are suited for key-value replacement. In this way, within the interval $t\in[\tau_{2},\tau_{3}]$, we gradually decrease the impact of the K-V AdaIN by combining the K-V replacements with a decreasing weight $\beta_{t}$ as follows:
@@
$\beta_{t}$ $=\frac{\beta_{\tau_{3}}-\beta_{\tau_{2}}}{\tau_{3}-\tau_{2}}\cdot(t-\tau_{2})+\beta_{\tau_{2}},$
in which, $\beta_{\tau_{2}}$ and $\beta_{\tau_{3}}$ are the predefined weighting factors.
By dynamically combining K-V AdaIN with K-V replacement in Eq. (11) and query blending in Eq. (9), as shown in Fig. 5(d), the stylization results achieve an effective balance between content fidelity and style richness. Further quantitative comparison with other stylization methods will typically be presented in Sec. 5.4.3.
@@ 4.3 Sliding-Window Consistent Smoothing
Flickering and artifacts in the final edited video occur when AdaIN-guided localized video stylization is applied frame by frame. We empirically observe that this issue is caused by the training-free stylization mechanism, which disrupts the temporal consistency of the edited latent. Inspired by previous works [1, 2], we further incorporate optical flow with specialized sliding-window consistent smoothing to improve inter-frame coherence during denoising.
Optical Flow Warping. Optical flow estimation is a pivotal technique widely employed in object tracking [39, 40]. It estimates pixel-wise horizontal and vertical motion velocities by analyzing pixel variations across frames. Building on this, we utilize the optical flow map to align semantic content across frames, enhancing the coherence of the edited video. Specifically, given two images, $A$ and $B$, we first employ RAFT [41] to estimate the optical flow map from $A$ to $B$. This map is then used to transform $B$, aligning it with $A$ to ensure precise spatial correspondence. In cases of occlusion, bidirectional optical flow can identify occluded regions and recover them with relevant content from $A$. The above process, termed $\operatorname{Warp}(A,B)$, is essential for the following discussions.
Sliding-Window Smoothing. Rather than relying on simple post‐processing, we integrate the Warp operation into the denoising process to achieve local smoothing. At time step $t$, we calculate the latent representation $\mathcal{Z}_{t\rightarrow 0}$ from the predicted noise $\varepsilon_{t}$ according to Eq. (2). $\mathcal{Z}_{t\rightarrow 0}$ is then decoded by the decoder $\mathcal{D}$ into a pixel representation $\{P_{t}^{i}\}_{i=1}^{N}$. For each decoded frame $P_{t}^{i}$, we apply a window of size $2m$ to warp neighboring frames onto it, followed by averaging to smooth the original pixel of $P_{t}^{i}$, as detailed below:
@@
The smoothed frame from the current window is used in the next window’s calculations, as shown in Fig. 2. Next, the VAE’s encoder $\mathcal{E}$ re-encodes the smoothed $\{\bar{P}_{t}^{i}\}_{i=1}^{N}$ into latent representations $\bar{\mathcal{Z}}_{t\rightarrow 0}$, which is a refined version of $\mathcal{Z}_{t\rightarrow 0}$. Subsequently, the predicted noise $\varepsilon_{t}$ is adjusted using the inversion of Eq. (2):
@@
Finally, the latent representation is updated according to the DDIM schedule in Eq. (1), enhancing coherence across frames:
@@
The above sliding-window consistent smoothing scheme is performed only within the interval $t\in[\tau_{4},\tau_{5}]$. Moreover, due to the challenges of optical flow estimation for fast-moving objects, we avoid global smoothing for long-range spatial alignment, opting instead for local sliding-window for short-range spatial alignment.
TABLE I: Quantitative comparison on DAVTG-WikiArt, where * denotes non-diffusion methods. The bold and underline indicate the best and second-best.
Methods Overall transfer performance Local style preservation Temporal consistency Inference efficiency
@@
TABLE II: Quantitative comparison on DAVTG-Laion, where * denotes non-diffusion methods. The bold and underline indicate the best and second-best.
Methods Overall transfer performance Local style preservation Temporal consistency Inference efficiency
[…]
Diffutoon [12] 38.610 23.700 0.563 0.959 0.953 0.979 44.263
@@ 5.1 Experimental Settings
Our proposed UniVST is based on the Stable Diffusion V-1.5 [32]. The hyper-parameters are configured as follows: $\tau_{0}=0.1T$, $\tau_{1}=0.2T$, $\tau_{2}=0.4T$, $\tau_{3}=1.0T$, $\tau_{4}=0.5T$, $\tau_{5}=0.6T$, $t_{0}=0.4T$, with the total time period $T$ set to 50. Additionally, we set $\gamma=0.35$, $\beta_{\tau_{2}}=0.1$, $\beta_{\tau_{3}}=0.9$, $r=0.3$, $k=15$, $m=2$, and $n=9$. Original videos are resized to $512\times 512$ pixels and processed in batches of 16 frames. We use the RAFT model [41] for optical flow estimation during the denoising process. UniVST can be executed on a single NVIDIA GeForce RTX 3090 GPU without additional training or fine-tuning, making it both efficient and practical for deployment.
@@
Datasets Details. Building upon the foundational work referenced in [17, 30], we have meticulously selected style images from two renowned databases: WikiArt [42] and Laion-Aesthetics-6.5+ [43], each contributing a collection of 20 distinct images. For the content videos, we have procured a dataset comprising 50 videos from the DAVIS2016 [44] and TGVE [45] datasets, which we combined and referred to as DAVTG. Through the integration of these datasets, we have successfully established two comprehensive subsets: DAVTG-WikiArt, which contains 1,000 curated pairs, and DAVTG-Laion, which also features 1,000 paired entries.
Baselines. We compare our UniVST with six advanced methods. These include two video stylization techniques: Diffutoon [12] and StyleCrafter [11]; two video editing frameworks: AnyV2V [13] and BIVDiff [10]; and two traditional non-diffusion methods: EFDM [20] and CAST [19]. Since our specific task does not perfectly align with some of the existing methods, we have made the necessary adjustments to ensure a fair comparison. In the case of Diffutoon [12], which is designed to process text inputs for coloring tasks and does not natively support the input of style images. To overcome this limitation, we leverage GPT-4o [46] to generate descriptive captions for the style images. For the style generation task with StyleCrafter [11], which does not inherently support video editing, we employ inversion to obtain the initial noise from the original video, which is then utilized in the generation process. For the unified video editing frameworks, AnyV2V [13] and BIVDiff [10], we have integrated InstantStyle [47] as the image style transfer model to perform the video style transfer task. Regarding the traditional non-diffusion methods, EFDM [20] and CAST [19], which are well-aligned with our task, we have retained them in their original form without any modifications.
5.2 Quantitative Comparison